# Robo-Instruct: Simulator-Augmented Instruction Alignment For Finetuning CodeLLMs

### v1

@@ Abstract
Large language models (LLMs) have shown great promise at generating robot programs from natural language given domain-specific robot application programming interfaces (APIs). However, the performance gap between proprietary LLMs and smaller open-weight LLMs remains wide. This raises a question: Can we fine-tune smaller open-weight LLMs for generating domain-specific robot programs to close the performance gap with proprietary LLMs? While Self-Instruct is a promising solution by generating a diverse set of training data, it cannot verify the correctness of these programs. In contrast, a robot simulator with a well-defined world can identify execution errors but limits the diversity of programs that it can verify. In this work, we introduce Robo-Instruct, which brings the best of both worlds — it promotes the diversity of Self-Instruct, while providing correctness of simulator-based checking. Robo-Instruct introduces RoboSim to synthesize a consistent world state on the fly by inferring properties relevant to the program being checked, and simulating actions accordingly. Furthermore, the instructions and programs generated by Self-Instruct may be subtly inconsistent — such as the program missing a step implied by the instruction. Robo-Instruct further addresses this with InstAlign, an instruction-program alignment procedure that revises the task instruction to reflect actual results of the generated program. Given a few seed task descriptions and the robot APIs, Robo-Instruct is capable of generating a training dataset using only a small open-weight model. This dataset is then used to fine-tune small open-weight language models, enabling them to even exceed the performance of several proprietary LLMs including GPT-3.5-Turbo and Gemini-Pro.
Figure 1: High-Level Overview of Robo-Instruct. This figure also illustrates an example of an invalid Self-Instruct-generated instruction and program, as well as pass@1 results of different LLMs on RoboEval.
@@ 1 Introduction
Large language models (LLMs) have demonstrated great promise at generating robot programs from natural language instructions [10, 12, 3, 18, 39, 17, 31, 11]. For example, consider an instruction for a service mobile robot: "Check how many conference rooms have no markers." The robot may be equipped with a domain-specific robot application programming interface (API) that includes skills such as go_to(location) for navigation and is_in_room(object) for perception. Since such domain-specific APIs do not exist in the training dataset of general-purpose LLMs, in-context learning (ICL) via few-shot examples is often employed to describe and use such APIs for performing few-shot inference. However, there is a significant performance gap [10] in the correctness of programs generated by ICL for large proprietary models and smaller open-weight models that can be deployed locally on robots. This raises a question: can we fine-tune small open-weight LLMs for generating domain-specific robot programs to close the performance gap with proprietary LLMs?
Since training datasets of the domain-specific robot programs are often unavailable, Self-Instruct might seem like a promising solution [36, 29]. Consider the setting of generating programs for service mobile robots that can perceive objects, navigate to various locations, manipulate items, and communicate with humans. By formulating these robot skills into APIs, we can create a few seed task examples demonstrating their use case and employ Self-Instruct to generate a diverse set of instruction-program pairs as training data, as illustrated in Fig. 1. However, using Self-Instruct naïvely may generate infeasible instructions—e.g., asking the robot to pick up multiple objects at once when it cannot due to physical constraints. They can also violate domain-specific constraints. For example, in Fig. 1, after line 2 confirms the absence of a key at the current location, line 3 erroneously attempts to pick up a key. Further, these instructions may not align with the generated programs, even if these programs are valid. For example, Fig. 1 shows an example instruction directing the robot to verbally ask in each room if a key exists, whereas the program instructs the robot to visually check in each room. Finally, the generated programs may have execution errors. These challenges may appear to be solvable using a simulator, but a simulator needs an initial world state to check against programs. A simulator using a hand-curated world state will end up rejecting the wide diversity of programs generated by Self-Instruct, even if they are executable, just because the world state did not capture some aspect relevant to them (e.g., the presence of a “key”).
@@
This work introduces Robo-Instruct, a new framework based on Self-Instruct, to address these issues and improve the performance of small open-weight language models for generating domain-specific robot programs. As shown in Fig. 1, Robo-Instruct introduces two novel components: (1) RoboSim, a task-agnostic simulator that encodes domain-specific constraints and validates robot programs generated from Self-Instruct. Critically, RoboSim dynamically synthesizes a consistent world state starting from arbitrary programs. (2) InstAlign, an instruction-program alignment procedure that revises the generated instructions to better reflect the intent of the generated programs. Robo-Instruct also employs a rejection-sampling mechanism that rejects invalid programs detected by RoboSim and queries Self-Instruct for a new program corresponding to the same generated instruction.
We validate Robo-Instruct by fine-tuning Codellama-Python-7B [30] and evaluate on RoboEval, a domain-specific code generation benchmark for service mobile robots. We show that Robo-Instruct is capable of improving the performance of the Codellama model by using only a small open-weight model to generate the training dataset. Compared to the base Codellama-Python-7B model without fine-tuning, our Robo-Instruct fine-tuned models outperform by $28.75\%$ in average pass@1 scores; and, compared to Self-Instruct fine-tuned model, our model outperform by $13.75\%$.; and the best pass@1 of Robo-Instruct fine-tuned model achieves a 68.75% match, surpassing the performance of the proprietary GPT-3.5-Turbo and Gemini-1.0-Pro.
@@
Our main contributions are as follows:  [^1] Website coming soon: https://amrl.cs.utexas.edu/robo-instruct/
@@
    We introduce Robo-Instruct, a new framework for improving the code generation performance of small open-weight language models for domain-specific robot programs. This framework introduces two novel components, RoboSim and InstAlign.
@@
    We introduce a dynamic world synthesis and evaluation process for generating relevant world states for automated code checking for diverse, arbitrary tasks in RoboSim.
@@
    We introduce InstAlign, an instruction alignment procedure to refine instruction-code pairs to improve alignment between instructions and code generated by Self-Instruct.
@@
    We fine-tune a small open-weight model, Codellama-Python-7B [30], using Robo-Instruct, and improve its performance to outperform several CodeLLMs, including Deepseek-Coder-33B [8], and Starcoder2-15B [21] and two proprietary LLMs, GPT-3.5-Turbo [27] and Gemini-1.0-Pro [33] on the RoboEval benchmark.
@@ 2 Robo-Instruct
In this section, we present how Robo-Instruct generates training datasets of domain-specific robot programs. Alg. 1 shows a broad overview of the framework. To add an entry in the training dataset, Self-Instruct first generates an instruction-program pair, $(\mathcal{I},\mathcal{P})$, from the robot APIs and seed tasks, shown in Appendix A.4. Then, RoboSim dynamically synthesizes a consistent world state on the fly as it executes and validates $\mathcal{P}$. If $\mathcal{P}$ is invalid, Robo-Instruct employs a rejection-sampling method, which generates a new program $\mathcal{P}$ given the same $\mathcal{I}$ and evaluates the new $\mathcal{P}$ again. This process repeats until $\mathcal{P}$ becomes valid or a predefined maximum resampling limit is reached. If the limit is reached, the instruction might be invalid given the domain-specific APIs or too complex to generate a program, so the instruction-program pair is discarded. Finally, if $\mathcal{P}$ is valid, InstAlign takes in $(\mathcal{I},\mathcal{P})$ to revise $\mathcal{I}$ to better reflect the intent of $\mathcal{P}$ and the aligned instruction and program is saved to the training dataset. In the following subsections, we elaborate on the specific design of each component.
Algorithm 1 Robo-Instruct: Instruction-Program Generation
@@
1:  Initialize: $\mathcal{D}=\emptyset$triangleright Training dataset
2:  Initialize: $N$triangleright Training dataset size
3:  Initialize: $m$triangleright Maximum resampling limit
4:  while $\text{len}(\mathcal{D})<N$ do
5:     $\mathcal{I}\leftarrow\textsc{Self-Instruct}_{inst}(\mathcal{S})$
6:     $\mathcal{P}\leftarrow\textsc{Self-Instruct}_{code}(\mathcal{S},\mathcal{I})$
7:     for $i=1$ to $m$ do
@@
[…]
10:           $\mathcal{P}\leftarrow\textsc{Self-Instruct}_{code}(\mathcal{S},\mathcal{I})$triangleright Rejection-sampling
11:        else
12:           $\mathcal{I}_{\text{aligned}}\leftarrow\textsc{InstAlign}(\mathcal{S},\mathcal{I},\mathcal{P})$triangleright Align instruction with the program
13:           $\mathcal{D}\leftarrow(\mathcal{I}_{\text{aligned}},\mathcal{P})$
14:           break
15:        end if
16:     end for
17:  end while
18:  return  $\mathcal{D}$
2.1 RoboSim: A Task-Agnostic Simulator For Domain-Specific Programs
@@
3:  Initialize: $\mathcal{W}_{\text{init}}$triangleright An initial world state with or without predefined information
@@
7:    $\mathcal{W}\leftarrow\mathcal{W}_{\text{init}}$triangleright Initialize a new world state
@@
[…]
2:  for $l\in p$ do $\triangleright$ Loop through every literal in the precondition
3:   if $\textsc{CheckDefined}(\mathcal{W},l)==\text{undefined}$ then
4:    $W\leftarrow\textsc{GrowWorld}(l,\mathcal{W})$triangleright Instantiate the literal and grow $\mathcal{W}$ to include it
5:   end if
6:  end for
7:  $\text{retval},\mathcal{W}\leftarrow\textsc{ExecUpdate}(\text{api\_fn},\text{params},\mathcal{W})$triangleright Execute api_fn and update $\mathcal{W}$
8:  return  retval, $\mathcal{W}$
We present a principled approach to design RoboSim for validating domain-specific robot programs. Alg. 2 illustrates the high-level algorithm used to assess the correctness of a robot program. RoboSim employs the concept of world state to simulate the robot actions directed by a program, ensuring consistent and reliable evaluation. A world state is a symbolic representation of the environment in which the robot operates, and it keeps track of the high-level changes in the robot state and the surrounding environment as the robot performs actions in order. For example, consider a program instruction that commands a robot to check if a pen is nearby. The world state queries the stored information about the surrounding environment, identifies all objects at the robot’s current location, and informs the program whether a pen is present.
However, since Self-Instruct generates arbitrary programs based on the provided APIs, RoboSim does not know what a plausible world state relevant to the program would be a priori — e.g., reasoning about the existence of a pen in the example program shown in Fig. 1. Thus, we equip RoboSim with the ability to synthesize the world state as more robot actions are performed. Our approach is inspired by angelic execution [4], which has previously been used for software verification of programs with partially defined library functions. In our case, instead of partially defined library functions, we have unknown plausible world states. RoboSim dynamically synthesizes and grows a world state based on domain-specific constraints (e.g., object permanence, robot skills, etc.) and the execution trace of the program, which allows it to infer a consistent and relevant world state.
Specifically, RoboSim modifies the program to replace all API calls with the DynamicEval function (Alg. 2 line 4) — when an API function is called during execution, the DynamicEval function is invoked instead.
DynamicEval makes an important extension to the formulation of STRIPS [7] to integrate with API functions. DynamicEval equips each API function with specific pre-conditions, effects, and return values. The pre-conditions are composed of literals tailored to the function’s requirements. For instance, the API function is_in_room("pen"), which determines if an object ‘pen’ is in the same room as the robot, uses two literals for its pre-condition: robot_at(X) and obj_at(X, "pen"). Generally, STRIPS assigns one of two possible values to each literal: True if the literal is defined, otherwise False. However, prior to program execution, DynamicEval is unaware of the program-relevant literals. Thus we assign a third value, undefined, to such unknown literals. Literals must thus be explicitly defined as either True or False, or they remain undefined if not specified.
Alg. 3 demonstrates how DynamicEval executes an API function and updates the world state. First, it calculates the precondition specified for the function. It then checks each literal in the precondition to see if it is defined. If a literal is undefined, DynamicEval invokes GrowWorld, a stochastic function that assigns a random truth value to the literal and updates the world state accordingly. Finally, DynamicEval proceeds to execute the API function using the current world state, retrieves the return values, and applies the function’s effects to update the world state.
Fig. 2 illustrates an example of RoboSim executing a generated program. Initially, RoboSim’s world state only specifies the robot’s current location, and whether a pen is in the same room as the robot remains undefined (line 2). Therefore, DynamicEval invokes GrowWorld to randomly determine a truth value for the obj_at(start_loc, "pen") literal, leading to two distinct execution paths depicted in light purple and blue. Subsequently, as additional API functions are called, more literals are introduced or updated in the world state to ensure consistent evaluations. In addition, a crucial aspect of RoboSim’s design is that a literal can also be set to an undefined value. In world state 2 on the purple execution path, after executing the pick("pen") function, the program is unable to determine if another pen exists in the current room, given the provided APIs. As a result, the world state updates the literal obj_at(start_loc, "pen") from True to Undefined.
Finally, due to the stochastic nature of DynamicEval, RoboSim must execute the generated program multiple times to validate the program. If all executions are successful, the program is deemed correct (Alg. 2 line 5-11).
Figure 2: Example of RoboSim executing a generated program and updating the world state. Initially, RoboSim begins with a world state that includes only the robot’s current location. As the program executes, two distinct execution paths emerge, depicted in light purple and blue. This figure demonstrates how the world state is updated along each execution path.
2.2 InstAlign: Instruction-Program Alignment Procedure
Given that LLMs are extensively trained in code understanding [30], InstAlign is a procedure that prompts an LLM to revise $\mathcal{I}$ to better reflect the intent of $\mathcal{P}$. This procedure involves two steps: first, given $\mathcal{I}$ and $\mathcal{P}$, InstAlign leverages Chain-of-Thought reasoning [37] (CoT) to prompt an LLM to generate a revised instruction, $\mathcal{I}_{\text{revised}}$; second, InstAlign invokes the LLM again to determine whether $\mathcal{I}$ or $\mathcal{I}_{\text{revised}}$ is more aligned with $\mathcal{P}$’s intent and output the chosen instruction as $\mathcal{I}_{\text{aligned}}$.
To generate $\mathcal{I}_{\text{revised}}$, the prompt to the LLM comprises the robot API function definitions, $\mathcal{I}$, $\mathcal{P}$, and CoT instructions. The CoT asks the LLM to perform the following three steps in order: 1. write down all the robot APIs used in the program; 2. examine these APIs and write down step by step what the program does; 3. combine all the information above to revise the robot instruction. Similarly, to determine $\mathcal{I}_{\text{aligned}}$, an LLM is prompted to think step by step about $\mathcal{P}$, $\mathcal{I}$ and $\mathcal{I}_{\text{revised}}$ to arrive at a conclusion. Detailed prompt is shown in Appendix A.6.

### v2

@@ Abstract
Open-weight LLMs are particularly appealing choices to generate training data for fine-tuning Code LLMs on domain-specific service robot applications because they are cost-effective, customizable, and offer better privacy protection. However, unlike proprietary LLMs, open-weight models are more error-prone and often produce programs that violate domain-specific constraints. A promising solution is to incorporate a robot simulator with a well-defined environment to verify program correctness. Yet, these environments require pre-enumeration of relevant entities and their states, which limits the diversity of programs that can be effectively verified. In this work, we introduce Robo-Instruct that preserves the diversity of programs generated by an LLM while providing the correctness of simulator-based checking. Robo-Instruct introduces RoboSim to dynamically synthesize consistent simulation environments for each generated program. Moreover, Robo-Instruct handles subtler instruction-program inconsistencies that do not result in a constraint violation via InstAlign, an LLM-aided instruction-program alignment process. Given domain-specific APIs and a few seed examples, Robo-Instruct can leverage an 8B Llama3 model to generate a training dataset for fine-tuning a 7B CodeLlama model. Our fine-tuned model achieves a 28.75% improvement in pass@1 over the original base model and a 13.75% improvement compared to its Self-Instruct-finetuned counterparts, even surpassing the performance of a few proprietary LLMs, such as GPT-3.5-Turbo and Gemini-Pro.
Figure 1: High-level overview of the Robo-Instruct framwork. This figure also shows the pass@1 score performance of Robo-Instruct fine-tuned LLM compared to other LLMs on RoboEval.
@@ 1 Introduction
Large language models (LLMs) have shown great promise at leveraging domain-specific application programming interfaces (APIs) and generating robot programs from natural language instructions (Hu et al., 2024; Huang et al., 2023b; Biggie et al., 2023; Liu et al., 2023a; Wu et al., 2023; Liang et al., 2022; Singh et al., 2023; Huang et al., 2023a). For instance, by formulating a robot’s navigation and perception skills into APIs, such as go_to(location) and is_in_room(object), an LLM can generate a program for a service mobile robot to complete the task: “Determine the number of conference rooms without markers”. However, despite impressive results, the performance gap remains wide between proprietary and open-weight LLMs in generating robot programs from domain-specific APIs (Hu et al., 2024). To bridge the performance gap, Self-Instruct (Wang et al., 2022) is a popular method for generating domain-specific data for finetuning LLMs. Given the definition of APIs and a few seed task examples, Self-Instruct prompts an LLM to generate diverse instruction-program pairs as training data.
Figure 2: Examples of programs violating domain-specific constraints.
This makes Self-Instruct a seemingly appealing way to prompt open-weight LLMs to generate training data, for the fine-tuning of Code LLMs on domain-specific service robot applications (due to their cost-effectiveness, better privacy protection, and customizability). However, open-weight LLMs are still prone to errors. As a result, using Self-Instruct naively can produce low-quality data, such as programs that violate domain-specific constraints or instructions that are infeasible for the robot to execute. For instance, as shown in the figure Fig. 2, the program may attempt to pick up an apple that is not present in the current location (example 1); or the instruction might ask the robot to pick up multiple objects simultaneously, which is physically impossible (example 2). To catch these errors, executing the program in a simulator with well-defined environments may appear promising. However, such environments require pre-enumeration of relevant entities and their states, which can result in rejecting a vast number of programs generated by Self-Instruct even if they are executable, simply because the pre-defined simulation environment fails to capture certain relevant aspects (e.g., the presence of an “apple”); this undermines the diversity of the data generated.
@@
This work introduces Robo-Instruct to bring the best of both worlds — it preserves the diversity of programs generated by an LLM while providing the correctness of simulator-based checking. Robo-Instruct is a new framework based on Self-Instruct, to address the challenges posed by using an open-weight LLM for generating domain-specific robot programs. As shown in Fig. 1, Robo-Instruct introduces two novel components: (1) RoboSim, a task-agnostic simulator that encodes domain-specific constraints and validates robot programs generated from Self-Instruct. Critically, RoboSim dynamically synthesizes consistent simulation environments starting from arbitrary programs. (2) InstAlign, an LLM-aided instruction-program alignment procedure that revises the generated instructions to better reflect the intent of the generated programs. Robo-Instruct also uses a rejection-sampling mechanism to discard invalid programs detected by RoboSim and query Self-Instruct for a new program based on the same instruction.
We use an 8B Llama3 model (AI, 2024) to generate instruction-program training data and fine-tune a 7B CodeLlama model (Rozière et al., 2024) on RoboEval, a domain-specific benchmark for service robot programs. Our Robo-Instruct-fine-tuned model significantly outperforms the base model by 28.75% in pass@1 scores and achieves a 13.75% improvement over its Self-Instruct-fine-tuned variant. Moreover, it surpasses several larger code models, including Deepseek-Coder-33B (Guo et al., 2024), Starcoder2-15B (Lozhkov et al., 2024), GPT-3.5-Turbo (OpenAI, 2022), and Gemini-Pro (Team et al., 2024). Finally, in real-world deployment, we demonstrate that Robo-Instruct achieves significantly lower latency compared with GPT-3.5 and GPT-4.
@@
Our main contributions are as follows  [^1] Project Website: https://amrl.cs.utexas.edu/robo-instruct/
@@
    We introduce Robo-Instruct, a new framework for improving open-weight LLMs’ ability to generate training data for finetuning Code LLMs on domain-specific service robot applications. This framework introduces two novel components, RoboSim and InstAlign.
@@
    We present RoboSim, a method to synthesize consistent simulation environments dynamically for verifying the generated programs. This method preserves the diversity of programs generated by an LLM while providing correctness with simulator-based checking.
@@
    We present InstAlign, a simple instruction-program alignment procedure that revises generated instructions better to reflect the actual results of the generated program.
@@
    We show that the Robo-Instruct finetuned model can significantly outperform the original base model and its finetuned variant using Self-Instruct in generating domain-specific service robot programs. It also surpasses several other larger Code LLMs.
  5. 5.
    We demonstrate that the Robo-Instruct fine-tuned model has significantly faster inference than proprietary models, making it more suitable for real-world robot deployments.
@@ 2 Robo-Instruct
In this section, we describe how Robo-Instruct generates training datasets for domain-specific service robot programs. Fig. 1 provides a high-level overview of the framework. To add an entry to the training dataset, Self-Instruct first generates an instruction-program pair based on robot APIs and seed tasks (detailed prompts in Appendix A.4.2). Then, RoboSim dynamically synthesizes a consistent simulation environment to validate the program as it executes.
If the program is invalid, Robo-Instruct applies a rejection-sampling method, which generates a new program based on the same instruction and re-evaluates it. This process continues until the program is valid or a predefined maximum resampling limit is reached. If the limit is exceeded, the instruction may be incompatible with the domain-specific APIs or too complex, so the instruction-program pair is discarded.
Finally, if the program is valid, InstAlign uses an open-weight LLM to refine the instruction to better align with the program’s intent, and the revised instruction-program pair is added to the training dataset. In the following subsections, we delve into the specific design of each component.
2.1 RoboSim: Dynamic Synthesis of Simulation Environments for Program Validation
We present a principled approach to designing RoboSim, a system for dynamically synthesizing consistent simulation environments to validate domain-specific robot programs. For service mobile robots, a simulation environment often relies on three concepts:
  1. 1.
    A list of entities to reason about, e.g., “apple”, “kitchen”
  2. 2.
    The type of the entities, and hence their affordances, e.g., “apple” is an object, you can pick it up; “kitchen” is a location, you can go to it, and it contains objects.
  3. 3.
    The state of the entities in the world, e.g., the “apple” is in the “kitchen”.
These concepts are closely related to the domain-specific APIs, where each API invocation during program execution can trigger updates to the simulation environment. To handle the interaction between APIs and simulation environments, we introduce DynamicEval, an algorithm inspired by Angelic Execution (Broy and Wirsing, 1981), a software engineering technique to infer program properties from incomplete API specifications. DynamicEval automatically generates a simulation environment for each program and checks its correctness within the inferred environment.
@@ 2.1 RoboSim: Dynamic Synthesis of Simulation Environments for Program Validation
1:  Input: api_fn $\triangleright$ The API function name
2:  Input: api_inputs $\triangleright$ The input received by the API invocation
3:  Input: $\mathcal{W}$triangleright The current simulation environment
4:  info $\leftarrow$ Extract_API_Invocation_Info(api_fn, api_inputs)
5:  entities $\leftarrow$ Infer_Relevant_Entities(info)
6:  for entity $\in$ entities do
7:     if Is_Entity_Initialized(entity, $\mathcal{W}$then
8:        entity_type $\leftarrow$ Deduce_Type(info)
9:        entity_new_state $\leftarrow$ Deduce_State(info)
@@
11:           Update_Entity_State(entity, info, $\mathcal{W}$)
12:        else
13:           raise “Error: state inconsistent or type mismatch”
14:        end if
15:     else
16:        Initialize_Entity_With_Random_State(entity, info, $\mathcal{W}$)
17:     end if
18:  end for
19:  retval $\leftarrow$ Get_Return_Value(info, $\mathcal{W}$)
20:  return  retval
As shown in Alg. 1, upon each API invocation, the corresponding inputs and the current simulation environment are passed into DynamicEval. DynamicEval first infers relevant entities when they appear in the program being checked (line 5). For instance, if a program includes the statement pick(”apple”), DynamicEval infers that apple is an entity to consider, even if it is not currently defined in the simulation environment.
For each relevant entity, if it has already been initialized (line 7), DynamicEval infers its type and new state from the API invocation and checks for any inconsistencies with the current simulation environment (lines 8-10). For example, pick requires an object type, while go_to requires a location type  [^2] In this example, type compatibility check is strict (i.e., “apple” is only an object and no further inference is made about its location). Nevertheless, the algorithm is also capable of handling more advanced scenarios.. If a program contains:
1def task_program():
2 pick("apple")
3 go_to("apple")
By invoking API calls sequentially, DynamicEval first infers that “apple” is an object and then raises an error when go_to(“apple”) is called. If no inconsistency is detected, the simulation environment is updated accordingly based on the API definition (line 11).
On the other hand, if the entity has not been initialized, it will be assigned a random plausible state (line 16). For example, the API is_in_room(object) checks if an object is in the same location as the robot and expects a boolean return. In this case, DynamicEval assigns a 50% probability for the apple to be in the robot’s current location — determining the state of the apple as either present or absent in the robot’s location:
1def task_program():
2 if is_in_room("apple"):
3 ...
Finally, DynamicEval computes and returns the value based on the API specification and the updated simulation environment. Fig. 3 illustrates how RoboSim leverages DynamicEval to dynamically synthesize the simulation environment and validate the program. In RoboSim, the simulation starts with only the robot at its initial position, and entities are added as the program runs. At execution timestep 2a, after the robot picks up an apple, it becomes unclear whether another apple remains at the location. Therefore, DynamicEval sets the apple’s state to ”Undefined” and removes it in subsequent executions. This method is also related to STRIPS planning, as we demonstrate the connection in Appendix A.2.
Figure 3: Illustration of RoboSim executing a task program and dynamically updating the simulation environment. Initially, the environment only contains the robot’s starting location. As the program runs, two possible execution paths emerge (purple and blue). The environment is updated by DynamicEval at each step, reflecting the robot’s actions, object states, and location changes along both paths.
In the example program shown in Fig. 3, it’s clear to humans that the program’s logic is flawed because it attempts to pick up the apple, which is not present in the room. But how would the simulator identify this as a failing state? RoboSim solves this issue by simulating all possible states of the discovered entities and verifying that none lead to erroneous program execution. In this case, the “apple” can either be present in the room or not. If the apple is not present, executing pick(”apple”) will result in an error. Checking all possible states requires exploring an exponentially growing number of combinations based on the entities discovered. To manage this, RoboSim employs a bounded compute budget to randomly sample from this exponential space, as detailed in the overall RoboSim algorithm in Alg. 2.
@@
3:  Initialize: $\mathcal{W}_{\text{init}}$triangleright An initial simulation environment with or without predefined information
@@
7:    $\mathcal{W}\leftarrow\mathcal{W}_{\text{init}}$triangleright Initialize a new simulation environment
@@
2.2 InstAlign: LLM-aided Instruction-Program Alignment Procedure
Figure 4: Overview of InstAlign.
Even after RoboSim verifies that a program doesn’t violate domain-specific constraints, it may still have subtle inconsistencies with the instructions — such as omitting a step implied by the instruction. The proposed rejection-sampling strategy alone cannot resolve this issue, and correcting verified programs to fully reflect the instructions while ensuring they remain valid is a challenging task.
We note that LLMs nowadays have demonstrated impressive code understanding capabilities (Rozière et al., 2024; Nam et al., 2024; Leinonen et al., 2023; Li et al., 2023; Lekshmi-Narayanan et al., 2024). Instead of correcting the program, it may be more effective for the LLM to revise its generated instructions to better align with the program’s intent.
We propose InstAlign, a procedure that prompts an LLM to revise generated instructions to better match the intent of the program, as shown in Fig. 4. InstAlign follows two steps: first, it uses Chain-of-Thought reasoning (CoT) (Wei et al., 2022) to generate a revised instruction based on the instruction-program pair; then, it prompts the LLM to compare the original and revised instructions, selecting the one that best aligns with the program’s intent. Detailed prompts are provided in Appendix A.4.3.